Hemophilia is an x-linked recessive trait. In a population of 2000 people, 4 women have hemophilia, 30 men have hemophilia, and 80 are carriers of the disorder. What is the frequency of the hemophilia allele? (Assume the population is 50% men and 50% women.)
a.0.059
b.0.033
c.0.057
d.0.039
e.0.038

In a population of 15,000 individuals, 1350 express a recessive trait. How many heterozygotes are there expected to be?
a.2100
b.3150
c.6300
d.1800
e.4200
Answer #1

Answer:

Q1). d. 0.039

Explanation:

4 hemophilic women = 8 hemophilic alleles

30 hemophilic men = 30 hemophilic alleles

80 carrieres = 80 normal alleles and 80 hemophilic alleles

Remaining individuals are normal = 2000 – 114 = 1886

943 men = 943 normal alleles

943 women = 943*2 = 1886 normal alleles

Total alleles = 3027

Total hemophilic alleles = 8+30+80 = 118

Frequency of hemophilic allele = 118/3027 = 0.039

Q2).

c. 6300

Explanation:

Recessive trait frequency = q^2 = 1350 / 15000 = 0.09

Recessive allele frequency = q = SQRT of q^2 = 0.3

Dominant allele frequency = p = 1-0.3 = 0.7

Heterozygote frequency = 2pq = 2*0.7 * 0.3 = 0.42

Number of heterozygotes = 0.42 * 15000 = 6300